OpenStudy (anonymous):

how to find lim x->0 (cos(1/x))/(1+(1/x))

OpenStudy (anonymous):

-1=<cos(1/x)<=+1 and lim (1/(1+1/x)=0 then 0*any value =0

OpenStudy (dumbcow):

change variable u = 1/x \[\lim_{u \rightarrow \infty} \frac{\cos u}{1+u}\] \[-1<\cos u<1\] thus \[\lim_{u \rightarrow \infty} \frac{\cos u}{1+u} = \frac{n}{\infty} = 0\]

OpenStudy (anonymous):

what does the -1<cosu<1 mean?

OpenStudy (anonymous):

@me100mks \(\cos x\) for any real \(x\) always gives a value between \(-1\) and \(1\). it should read \(\forall x\in\mathbb{R},-1\le \cos x\le1\)